com2sec secName  ipSource community

default

group  groupName  model  secName

v1
v2c
usm

access  groupName  context  model  level  prefix readScope  writeScope notifyScope
""        v1          noauth  exact  viewName viewName viewName

          v2c         auth    prefix  none        none        none
          usm         priv

          any

view  viewName  type  oidSubtree [mask]

included
excluded